# Decoherence in Exchange-Coupled Quantum Spin Qubit Systems: Impact of Multiqubit Interactions and Geometric Connectivity

### v1

@@ Abstract
We investigate the impact of different connectivities on the decoherence time in quantum systems under quasi-static Heisenberg noise. We considered three types of fundamental units, including node, stick and triangle and connect them into rings, chains, and trees. We find that rings exhibit greater stability compared to chains, contrary to the expectation that higher average connectivity leads to decreased stability. Additionally, the “stick” configuration is more stable than the “triangle” configuration. We also observe similar trends in entanglement entropy and return probability, indicating their potential use in characterizing decoherence time. Our findings provide insights into the interplay between connectivity and stability in quantum systems, with implications for the design of robust quantum technologies and quantum error correction strategies.
@@ I Introduction
To improve the fidelity of multiqubit systems, it is important to investigate the factors that may impact their decoherence, and subsequently, the fidelity of applied quantum controls as the system size increases Reagor et al. (2018); Fang et al. (2023); Kim et al. (2022). In order to execute quantum algorithms, qubits are necessarily connected together with possible operations on the connected ones. However, such connectivity usually implies decoherence, as can be seen in the case of ultracold atom systems and IBM’s superconducting qubits Bloch (2008); Zhang et al. (2017). In a quantum device, two aspects on how the qubits are connected worth consideration: the number of links between the pairs of qubits, as well as how the qubits are geometrically connected. Intuitively, if there are less connections between qubits, the device would have longer coherence time because the qubits are less affected by their neighbors. However, as shown in an important prior work Throckmorton and Das Sarma (2022), such perception is not entirely accurate and the geometry of the device, such as ring and chain configuration, is also important. It has been shown that larger average connectivity does not necessarily lead to faster decay of coherence. Nevertheless, the effect of geometric connectivity on decoherence time remains an open question. Therefore, in this work, our aim is to investigate more comprehensively into this aspect and provide a more thorough understanding on how the geometric connectivity affects decoherence in multiqubit systems.
@@ II Model and method
where $\bm{\sigma}_{i}$ is the vector of Pauli matrices on spin $i$. Experimental evidence suggests that the Heisenberg interaction is the dominant source of noise between qubits, denoted by $J_{ij}(t)$. By studying the evolution of decoherence time under different connectivity configurations, we aim to gain a comprehensive understanding of the relationship between geometric connectivity and decoherence in multiqubit systems.
In the presence of Heisenberg interaction, it is important to consider both static and dynamic aspects of the noise. In experimental settings, the frequency of noise fluctuations is much lower than the characteristic timescale of the experimental manipulations. Therefore, we primarily focus on the effect of quasi-static noise, which is characterized by a normal distribution. To account for this noise, we employ the Monte Carlo method to randomly select noise intensities and average their effects. Specifically, for quasi-static noise, we model the coupling strength $J_{ij}(t)$ as a Gaussian distribution with a mean value $J_{ij,0}$ and a deviation $\sigma$. It is worth noting that as experimental operations become more accurate in the future, the influence of dynamic noise may become more significant. In this paper, our main focus is on the study of quasi-static noise and its impact. The effects of dynamic noise are discussed in Appendix A.
Next, we incorporate the study of geometric connectivity into our model. We consider several different elementary units, including the “node”, “stick”, and “triangle”, and explore the geometric connectivity configurations of ring, chain, and tree. As depicted in Fig. 1, the vertices represent qubits, the gray lines represent the parts within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. These elementary units are then geometrically connected in different configurations, in the forms of a ring, chain, or tree, as shown in Fig. 2. In this study, we investigate the decoherence time of these elementary units with $L$ ranging between 4 and 10, where $L$ represents the number of qubits.
@@
Figure 2: Illustration of chain, ring and tree configurations using different elementary units. The vertices represent qubits, the gray lines indicate the regions within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. The arrows inside the vertices represent the initial states that we choose. An upward arrow indicates a spin-up state at this site, while a downward arrow indicates a spin-down state. In our study, we focus on the qubits that are affected by noise-induced decoherence, which are the qubits between which quantum gates are applied. Moreover, only connected qubits can implement two-qubit gates. Therefore, we adopt a model with Heisenberg interaction between connected qubits to investigate the impact of noise.
@@
where $|\Psi(0)\rangle$ is the initial state and $|\Psi(t)\rangle$ is the state at time $t$. We choose the initial state $|\Psi(0)\rangle=|\uparrow\downarrow\cdots\rangle$, as shown in Fig. 2. The return probability $P(t)\propto P_{A}(t)P_{\phi}(t)$ can be expressed as
@@
where $P_{A}(t)$ is the amplitude term, $P_{\phi}(t)$ is the frequency term, and the decoherence time $T^{*}_{2}$ is extracted from fitting. In the case of elementary unit “node” with quasi-static noise, we analyze the probability $P(t)$ as a function of time $t$. Fig. 3 shows the evolution of $P(t)$ for $L=6,J_{0}=100/t_{0},\sigma=0.4/t_{0}$, represented by the solid line. In this work, we denote $t_{0}$ as our time unit, which is about 10 ns in superconducting qubits and about 1 $\mu s$ for a typical iron trap quantum device. The dashed line corresponds to the fitted curve, while the dotted line represents the asymptotic value $P_{\infty}$. We extract the envelope lifetime using the Eq. (3). The obtained decoherence time is $T_{2}^{*}/t_{0}=0.489$. This quantity is commonly measured in experimental settings and provides valuable insights into the system’s decoherence dynamics.
Figure 3: The probability $P(t)$ as a function of time $t$ for elementary unit “node”, with quasi-static noise under chain configuration. The solid line represents the evolution of $P(t)$ as a function of time under $L=4,J_{0}=100/t_{0},\sigma=0.4/t_{0}$, $t_{0}$ is the time unit, while the dashed line corresponds to the fitted curve. We extract the envelope lifetime based on Eq. (3) to characterize the decoherence time, and we obtain $T^{*}_{2}=0.706/t_{0}$.
@@ III.1 Decoherence time under quasi-static noise
We initially focus on the ring and chain connectivity configurations. We will discuss the tree configuration in a later stage because the system size involved may be significantly larger. We specifically investigate the results of the “node” elementary unit’s tree configuration for a system size of $L=8$ in Sec. III.2.
@@
Firstly, we extract the decoherence time $T_{2}^{*}$ for the elementary unit “node”, with the system size up to $L=10$. We set the interaction strength as $J_{0}=100/t_{0}$ and the deviation as $\sigma=0.5/t_{0}$. Choosing $J_{0}=100/t_{0}$ allows for faster decay of the system and more accurate calculation of $T_{2}^{*}$, as the product of the energy unit $J_{0}$ and the decoherence time remains constant. Fig. 4 shows that the decoherence time $T_{2}^{*}$ exhibits similar behavior to that reported in Ref. Throckmorton and Das Sarma (2022). It can be observed that in the ring configuration, with the “node” as the elementary unit, the overall decoherence time decreases as the system size $L$ increases. Specifically, the decoherence time decreases from $T_{2}^{*}=1.004/t_{0}$ at $L=4$ to $T_{2}^{*}=0.201/t_{0}$ at $L=10$, resulting in a decline of approximately $80\%$. Moreover, the even-qubit systems exhibit longer lifetimes compared to the odd-qubit systems. On the other hand, for the chain configuration, the decoherence time also decreases with increasing system size $L$, but at a slower rate compared to the ring configuration. Similar results have been reported in Ref. Throckmorton and Das Sarma (2022). Specifically, for the chain, the decoherence time decreases from $T_{2}^{*}=0.543/t_{0}$ at $L=4$ to $T_{2}^{*}=0.297/t_{0}$ at $L=10$, corresponding to a decline of approximately $45\%$. Minor discrepancies between our numerical results and those of Ref. Throckmorton and Das Sarma (2022) may arise from differences in the chosen energy scale, variations in the Monte Carlo process, and numerical fluctuations when extracting $T_{2}^{*}$ from the envelope fitting.
@@
Figure 4: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary unit “node”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decay trend is similar to Ref. Throckmorton and Das Sarma (2022).
@@
Next, we investigate the behavior of elementary units “stick” and “triangle” up to $L=10$, considering an interacting strength of $J_{0}=100/t_{0}$ and a deviation of $\sigma=0.5/t_{0}$. In Fig. 5, we observe that for the ring geometry with the “stick” as the elementary unit, the overall decoherence time decreases as the system size $L$ increases. Specifically, the decoherence time decreases by $48\%$, from $T_{2}^{*}=1.185/t_{0}$ at $L=6$ to $T_{2}^{*}=0.614/t_{0}$ at $L=10$. We note that we excluded the case of $L=4$ for the “stick” elementary unit, as it comprises only two elementary units, which cannot form a ring. As for the chain geometry with the “stick” as the elementary unit, the overall decoherence time also decreases with increasing system size $L$. Comparatively, at $L=6$, the decoherence time decreases from $T_{2}^{*}=0.625/t_{0}$ to $T_{2}^{*}=0.444/t_{0}$ at $L=10$, representing a decrease of $29\%$. In summary, the chain exhibits a slower rate of decrease in decoherence time with increasing size. However, the ring geometry consistently displays greater stability than the chain at the same system size.
@@
Figure 5: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary units “stick”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decoherence time for ring configuration is longer than that of chain configuration.
@@
In Fig. 6, we observe consistent trends in the decoherence time for both ring and chain geometries. For the ring geometry with the “triangle” as the elementary unit, the decoherence time decreases gradually as the system size $L$ increases. Specifically, $T_{2}^{*}$ decreases from $0.655/t_{0}$ for $L=4$ to $0.414/t_{0}$ for $L=10$. On the other hand, in the chain geometry with the “triangle” as the elementary unit, the overall decoherence time also decreases with increasing system size $L$, showing a similar rate of decrease. The decoherence time decreases from $T_{2}^{*}=0.462/t_{0}$ for $L=4$ with to $T_{2}^{*}=0.107/t_{0}$ for $L=10$. The overall rate of decay is relatively similar between the ring and chain geometries, highlighting the distinct influence of the elementary unit “triangle” and “stick” on the decoherence dynamics.
@@
Figure 6: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary units “triangle”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decoherence time for ring configuration is also longer than that of chain configuration.
@@
The average number of connections between qubits in a ring configuration consistently exceeds that found in a chain configuration. This observation leads to an intuitive hypothesis: a lower average number of qubit connections could potentially extend the decoherence time. The rationale behind this is straightforward — fewer qubit connections imply limited positions for the application of quantum gates, which in turn may reduce the impact of noise effects on the system. This relationship between the complexity of qubit interconnections and decoherence dynamics merits further exploration, as it could offer insights into optimizing quantum computing architectures for enhanced stability. However, in Fig. 5 and Fig. 6, we can conclude that the configuration ring is in general more stable than chain.
@@
(a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption
@@
In Figures 7(c) and 7(d), we investigate the decoherence times of the elementary unit “triangle” for both ring and chain configurations, taking into account varying values of deviation $\sigma$ and interaction strength $J_{0}=100/t_{0}$. Similar to the “stick” configuration, the ring geometry consistently shows longer overall decoherence times compared to the chain. Unlike the “stick” configuration, however, the ring demonstrates weaker dependence to system size for the elementary unit “triangle”. This implies that, irrespective of the $\sigma$ value, the decoherence time in the ring configuration decreases at a slower rate as the system size increases. Nonetheless, as observed in earlier comparisons, the ring configuration continues to outperform the chain in terms of stability, maintaining longer decoherence times than the chain at the same system size.
@@
Regarding the observation that rings are more stable than chains, we speculate that this behavior may be attributed to symmetry. For instance, in a three-qubit system without considering variance, the Hamiltonian of the ring commutes with the total spin Buterakos and Das Sarma (2021), thereby conserving the total spin and providing some protection to the quantum state against evolving into other subspaces. Although the Hamiltonian of the ring does not commute with the total spin in larger systems, high symmetry may still protect the quantum state and make it more challenging to leak into other subspaces.

### v2

@@ Abstract
We investigate the impact of different connectivities on the decoherence time in quantum systems under quasi-static Heisenberg noise. We considered three types of elementary units, including node, stick and triangle and connect them into ring, chain, and tree configurations. We find that rings exhibit greater stability compared to chains, contrary to the expectation that higher average connectivity leads to decreased stability. Additionally, the stick configuration is more stable than the triangle configuration. We also observe similar trends in entanglement entropy and return probability, indicating their potential use in characterizing decoherence time. Our findings provide insights into the interplay between connectivity and stability in quantum systems, with implications for the design of robust quantum technologies and quantum error correction strategies.
@@ I Introduction
To improve the fidelity of multiqubit systems, it is important to investigate the factors that may impact their decoherence, and consequently, the fidelity of applied quantum controls as the system size increases Reagor et al. (2018); Fang et al. (2023); Kim et al. (2022). For the execution of quantum algorithms, qubits must be interconnected, allowing for operations to be performed on the connected qubits. However, this connectivity often leads to decoherence, as observed in systems such as ultracold atom setups and IBM’s superconducting qubits Bloch (2008); Zhang et al. (2017). In a quantum device, two aspects on how the qubits are connected worth consideration: the number of links between the pairs of qubits, as well as how the qubits are geometrically connected. Intuitively, if there are less connections between qubits, the device would have longer coherence time because the qubits are less affected by their neighbors. Contrary to the common perception that increased connectivity invariably leads to faster decoherence, an important previous study Throckmorton and Das Sarma (2022) has demonstrated that this is not always the case. The research highlighted that the geometry of the device, such as ring and chain configurations, plays a significant role in determining decoherence behavior. It was found that greater average connectivity does not necessarily result in a more rapid decay of coherence. This revelation opens up a new area of inquiry regarding the impact of geometric connectivity on decoherence times. Inspired by this work, our research aims to provide a more comprehensive understanding of how geometric connectivity influences decoherence in multiqubit systems.
@@ II Model and method
where $\bm{\sigma}_{i}$ is the vector of Pauli matrices on spin $i$. Experimental evidence suggests that the Heisenberg interaction is the dominant source of noise between qubits, denoted by $J_{ij}(t)$, which is usually referred to as “crosstalk” in the literature Throckmorton and Das Sarma (2022). In our analysis, we have chosen to disregard single-qubit interaction terms (such as Zeeman terms represented by $E_{z}\sigma_{iz}$) to primarily focus on two-qubit crosstalk, because the errors incurred in two-qubit gates are much higher than those in single-qubit gates Johnstun and Van Huele (2021). In Appendix A, we have shown results that include a single-qubit term with $E_{Z}=0.1J_{0}$ (where $J_{0}$ is a parameter relevant to the strength of crosstalk to be described later), and have found that our main conclusions remain valid.
By examining the evolution of decoherence time across different connectivity configurations, we aim to achieve a comprehensive understanding of how geometric connectivity influences decoherence in multiqubit systems.
Our research focuses on the evolution of a given initial state $|\Psi(0)\rangle$ under the influence of such noise in Eq. (1). In other words, connected quantum bits experience crosstalk, which leads to decoherence of a prepared initial state. We investigate how long a prepared initial state can be preserved under this noisy situation in different geometric configurations, which can be understood as undergoing a “scrambling” process where a closed many-body system loses memory of what it was.
In the presence of Heisenberg interaction, it is important to consider both static and dynamic aspects of the noise. In experimental settings, the frequency of noise fluctuations is much lower than the characteristic timescale of the experimental manipulations. Therefore, we primarily focus on the effect of quasi-static noise, which is characterized by a normal distribution. To account for this noise, we employ the Monte Carlo method to randomly select noise intensities and average their effects. Specifically, for quasi-static noise, we model the coupling strength $J_{ij}(t)$ as a Gaussian distribution with a mean value $J_{ij,0}$ and a deviation $\sigma$. We use Monte Carlo method to realize the Gaussian distribution. For example, the interaction between qubits 1 and 2 is randomly sampled as $J_{12}\sim N(J_{0},\sigma^{2})$ and the interaction between qubits 2 and 3 is randomly sampled as $J_{23}\sim N(J_{0},\sigma^{2})$, etc. Note that although these $J_{ij}$ are sampled randomly, during one Monte Carlo step they remain unchanged, implying a quasi-static noise. It is worth noting that as experimental operations become more accurate in the future, the influence of dynamic noise may become more significant. In this paper, our main focus is on the study of quasi-static noise and its impact. The effects of dynamic noise are discussed in Appendix B.
Next, we expand our model to include the study of geometric connectivity. We explore various elementary units such as the “node”, “stick”, and “triangle,” and examine their geometry in ring, chain, and tree configurations to understand how these different connectivities affect the system’s behavior. As depicted in Fig. 1, the vertices represent qubits, the gray lines represent the parts within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. These elementary units are then geometrically connected in different configurations, in the forms of a ring, chain, or tree, as shown in Fig. 2. In this study, we investigate the decoherence time of these elementary units with $L$ ranging between 4 and 10, where $L$ represents the number of qubits. A visualization of these larger clusters are shown in Appendix C.
@@
Figure 2: Illustration of chain, ring and tree configurations using different elementary units. The vertices represent qubits, the gray lines indicate the regions within each unit where quantum gates are allowed, and the red lines represent the connections between repeated units. The arrows inside the vertices represent the initial states that we choose. An upward arrow indicates a spin-up state at this site, while a downward arrow indicates a spin-down state. In our study, we focus on the qubits that are affected by noise-induced decoherence, which are the qubits between which quantum gates are applied. Moreover, only connected qubits can implement two-qubit gates. Therefore, we adopt a model with Heisenberg interaction between connected qubits to investigate the impact of the crosstalk noise.
@@
where $|\Psi(0)\rangle$ is the initial state and $|\Psi(t)\rangle$ is the state at time $t$. We choose the initial state as $|\Psi(0)\rangle=|\uparrow\downarrow\cdots\rangle$, shown in Fig. 2. We note that the decoherence of a quantum system strongly depends on its initial state, as studied previously in Hung et al. (2013, 2014). In this work, our aim is to conduct a comparative analysis of various qubit geometric configurations, which necessitates the selection of an initial state that is compatible with these different setups. The product state selected here is the most basic form that can accommodate all configurations, ensuring a fair and unbiased comparison. We have also explored the GHZ state as an alternative initial state. The findings related to this are detailed in Appendix D, and they do not alter our primary conclusion.
The return probability $P(t)\propto P_{A}(t)P_{\phi}(t)$ can be expressed as
@@
where $P_{A}(t)$ is the amplitude term, $P_{\phi}(t)$ is the frequency term, and the decoherence time $T^{*}_{2}$ is extracted from fitting. In the case of elementary unit “node” with quasi-static noise, we analyze the probability $P(t)$ as a function of time $t$. Fig. 3 shows the evolution of $P(t)$ for $L=6,J_{0}=100/t_{0},\sigma=0.4/t_{0}$, represented by the solid line. In this work, we denote $t_{0}$ as our time unit, which is about 100 $\mu$s in superconducting qubits Rigetti et al. (2012) and about 700 $\mu$s for a typical iron trap quantum device Zhang et al. (2020). The dashed line corresponds to the fitted curve, while the dotted line represents the asymptotic value $P_{\infty}$. We extract the envelope lifetime using the Eq. (3), and obtain decoherence time is $T_{2}^{*}=0.489t_{0}$. In the case of a transmon qubit system, $T_{2}^{*}$ is around 95 $\mu$s, in consistency with experiments.
Figure 3: The probability $P(t)$ as a function of time $t$ for elementary unit “node”, with quasi-static noise under chain configuration. The solid line represents the evolution of $P(t)$ as a function of time under $L=4,J_{0}=100/t_{0},\sigma=0.4/t_{0}$, $t_{0}$ is the time unit, while the dashed line corresponds to the fitted curve. We extract the envelope lifetime based on Eq. (3) to characterize the decoherence time, and we obtain $T^{*}_{2}=0.489t_{0}$.
@@ III.1 Decoherence time under quasi-static noise
We initially focus on the ring and chain connectivity configurations. We will discuss the tree configuration in a later stage because the system size involved is significantly larger. We specifically investigate the results of the “node” elementary unit’s tree configuration for a system size of $L=8$ in Sec. III.2.
@@
Firstly, we extract the decoherence time $T_{2}^{*}$ for the elementary unit “node”, with the system size up to $L=10$. We set the interaction strength as $J_{0}=100/t_{0}$ and the deviation as $\sigma=0.5/t_{0}$. Choosing $J_{0}=100/t_{0}$ allows for faster decay of the system and more accurate calculation of $T_{2}^{*}$, as the product of the energy unit $J_{0}$ and the decoherence time remains constant. Fig. 4 shows that the decoherence time $T_{2}^{*}$ exhibits similar behavior to that reported in Ref. Throckmorton and Das Sarma (2022). We fit these results using the power law form $T_{2}^{*}(L)=\tau_{0}L^{-\gamma}$, finding the corresponding fitting parameters $(\tau_{0},\gamma)$ for ring are (4.957, 1.192), and for chain are (2.950, 1.197). It can be observed that in the ring configuration, with the “node” as the elementary unit, the overall decoherence time decreases as the system size $L$ increases. Specifically, the decoherence time decreases from $T_{2}^{*}=1.004t_{0}$ at $L=4$ to $T_{2}^{*}=0.201t_{0}$ at $L=10$, resulting in a decline of approximately $80\%$. On the other hand, for the chain configuration, the decoherence time also decreases with increasing system size $L$, but at a slower rate compared to the ring configuration. Similar results have been reported in Ref. Throckmorton and Das Sarma (2022). Specifically, for the chain, the decoherence time decreases from $T_{2}^{*}=0.543t_{0}$ at $L=4$ to $T_{2}^{*}=0.297t_{0}$ at $L=10$, corresponding to a decline of approximately $45\%$. Minor discrepancies between our numerical results and those of Ref. Throckmorton and Das Sarma (2022) may arise from differences in the chosen energy scale, variations in the Monte Carlo process, and numerical fluctuations when extracting $T_{2}^{*}$ from the envelope fitting.
In Fig. 4, we notice an oscillatory behavior in the dephasing time for the ring configuration of nodes. Specifically, there are noticeable dips for configurations with an odd number of qubits, suggesting that these configurations have a shorter lifetime compared to similar cases with an even number of qubits. This phenomenon can be attributed to an “even-odd” effect, which is linked to the number of spins in the system. When the number of spins is even, each spin can align itself in the opposite direction to its two neighbors. However, in a system with an odd number of spins, at least one spin will be unable to align oppositely to both its neighbors, resulting in a state of “frustration.” This frustrated state has a higher local entropy, making it less stable and leading to shorter lifetimes. This effect is specific to ring configurations with an odd number of nodes and does not occur in configurations based on sticks and triangles, where the number of qubits is always even. Similarly, chain configurations do not exhibit this frustration effect. In the analysis of GHZ states presented in Appendix D, we do not observe this even-odd effect, as expected, because GHZ states are symmetrized and thus not subject to frustration.
@@
Figure 4: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary unit “node”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The dashed line represents the fitting curve for the ring using the formula $T_{2}^{*}(L)=\tau_{0}L^{-\gamma}$, while the dash-dotted line represents the fitting curve for the chain. For the ring configuration, the fitting parameters are $(\tau_{0},\gamma)$ = (4.957, 1.192), and for the chain configuration, the parameters are (2.950, 1.197). The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decay trend is similar to Ref. Throckmorton and Das Sarma (2022).
@@
Next, we investigate the behavior of elementary units “stick” and “triangle” up to $L=10$, considering an interacting strength of $J_{0}=100/t_{0}$ and a deviation of $\sigma=0.5/t_{0}$. In Fig. 5, we observe that for the ring configuration with the “stick” as the elementary unit, the overall decoherence time decreases as the system size $L$ increases. Specifically, the decoherence time decreases by $48\%$, from $T_{2}^{*}=1.185t_{0}$ at $L=6$ to $T_{2}^{*}=0.614t_{0}$ at $L=10$. We note that we excluded the case of $L=4$ for the “stick” elementary unit, as it comprises only two elementary units, which cannot form a ring. We have fitted the results using the power law form $T_{2}^{*}(L)=\tau_{0}L^{-\gamma}$. The corresponding fitting parameters $(\tau_{0},\gamma)$ for the ring configuration are (12.008, 1.293), and for the chain configuration, they are (1.589, 0.606). As for the chain configuration with the “stick” as the elementary unit, the overall decoherence time also decreases with increasing system size $L$. Comparatively, at $L=6$, the decoherence time decreases from $T_{2}^{*}=0.625t_{0}$ to $T_{2}^{*}=0.444t_{0}$ at $L=10$, representing a decrease of $29\%$. In summary, the chain exhibits a slower rate of decrease in decoherence time with increasing size. However, the ring configuration consistently displays greater stability than the chain at the same system size.
@@
Figure 5: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary units “stick”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain configuration. The dashed line represents the fitting curve for the ring using the formula $T_{2}^{*}(L)=\tau_{0}L^{-\gamma}$, while the dash-dotted line represents the fitting curve for the chain. For the ring configuration, the fitting parameters are $(\tau_{0},\gamma)$ = (12.008, 1.293), and for the chain configuration, the parameters are (1.589, 0.606). The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decoherence time for ring configuration is longer than that of chain configuration.
@@
In Fig. 6, we observe consistent trends in the decoherence time for both ring and chain configurations. For the ring configuration with the “triangle” as the elementary unit, the decoherence time decreases gradually as the system size $L$ increases. Specifically, $T_{2}^{*}$ decreases from $0.655t_{0}$ for $L=4$ to $0.414t_{0}$ for $L=10$. On the other hand, in the chain configuration with the “triangle” as the elementary unit, the overall decoherence time also decreases with increasing system size $L$, showing a similar rate of decrease. The decoherence time decreases from $T_{2}^{*}=0.462t_{0}$ for $L=4$ with to $T_{2}^{*}=0.107t_{0}$ for $L=10$. Similarly, the fitting parameters $(\tau_{0},\gamma)$ for the ring configuration are (1.363, 0.511), , and for the chain configuration, they are (3.182, 1.371) when using the power law form $T_{2}^{*}(L)=\tau_{0}L^{-\gamma}$. The overall rate of decay is relatively similar between the ring and chain configurations, highlighting the distinct influence of the elementary unit “triangle” and “stick” on the decoherence dynamics.
@@
Figure 6: The decoherence time $T_{2}^{*}$ as a function of system size $L$ for elementary units “triangle”, with quasi-static noise. The blue inverted triangle markers correspond to the ring configuration, while the orange upright triangle markers represent the chain The dashed line represents the fitting curve for the ring using the formula $T_{2}^{*}(L)=\tau_{0}L^{-\gamma}$, while the dash-dotted line represents the fitting curve for the chain. For the ring configuration, the fitting parameters are $(\tau_{0},\gamma)$ = (1.363, 0.511), and for the chain configuration, the parameters are (3.182, 1.371). The interacting strength $J_{0}=100/t_{0}$ and the deviation $\sigma=0.5/t_{0}$. The decoherence time for ring configuration is also longer than that of chain configuration.
@@
Intuitively, one might expect that a lower average number of qubit connections would extend the decoherence time. The reasoning is straightforward: fewer qubit connections mean fewer positions for applying quantum gates, which could potentially reduce the impact of noise on the system. However, in Fig. 5 and Fig. 6, we observe that the ring configuration is generally more stable than the chain configuration, despite the fact that the average number of connections between qubits in a ring consistently exceeds that in a chain. This counterintuitive relationship between the complexity of qubit interconnections and decoherence dynamics could provide valuable insights into optimizing quantum computing architectures for improved stability. We also observe that the decoherence times for configurations formed by elementary units “stick“ and “triangle“ do not exhibit the oscillatory dependence seen in Fig. 4. This absence of oscillatory behavior is due to the fact that these configurations always contain an even number of qubits, eliminating the “even-odd” effect.
@@
(a) Refer to caption (b) Refer to caption
(c) Refer to caption (d) Refer to caption
@@
In Figures 7(c) and 7(d), we investigate the decoherence times of the elementary unit “triangle” for both ring and chain configurations, taking into account varying values of deviation $\sigma$ and interaction strength $J_{0}=100/t_{0}$. Similar to the “stick” configuration, the ring configuration consistently shows longer overall decoherence times compared to the chain. Unlike the “stick” configuration, however, the ring demonstrates weaker dependence to system size for the elementary unit “triangle”. This implies that, irrespective of the $\sigma$ value, the decoherence time in the ring configuration decreases at a slower rate as the system size increases. Nonetheless, as observed in earlier comparisons, the ring configuration continues to outperform the chain in terms of stability, maintaining longer decoherence times than the chain at the same system size.
@@
The stability of different qubit configurations under noise can be influenced by various factors, but we propose that symmetry plays a crucial role in the superior stability of ring configurations compared to chains. Taking a small system with $L=3$ as an example Buterakos and Das Sarma (2021), the Hamiltonian for the ring configuration commutes with the total spin, meaning that the total spin is conserved. This conservation provides a degree of protection against decoherence, preserving the quantum state. Chain and tree configurations do not exhibit this property. As the size of the system increases, the total spin conservation is no longer maintained in the ring configuration. However, a residual protective effect from the initial symmetry persists. We attribute this lingering protection to the inherent symmetry of the ring configuration, which we believe is the primary reason for its enhanced stability compared to chains.